OpenStudy (anonymous):

Can someone please HELP me solve this problem (e^t/3 /2^t/2)^6 -2^-3t e^4 =0

OpenStudy (anonymous):

it could be simplify (if I read it correctly) as: e^(2t)/2^(3t) -2^(-3t) *e^4=0 multiply both sides by 2^(3t), then: e^(2t) - e^4=0 e^(2t) = e^4 which means that 2t=4 t=2